OpenStudy (kkbrookly):

Determine whether each sequence converges or diverges If it converges, give the limit. 1) an=n-3/5n+2

OpenStudy (anonymous):

the horizontal asymptote of \[f(x)=\frac{x-3}{5x+2}\] is \[y=\frac{1}{5}\]

OpenStudy (anonymous):

it is the same question, only stated differently and in a later class probably, you did this in a precalc class i am sure it is the same thing exactly

OpenStudy (anonymous):

assuming of course that it was \[\lim_{n\to \infty}\frac{n-3}{5n+2}\]

OpenStudy (kkbrookly):

I am doing this in precal, but this is the first day I am working on it and I have absolutely no clue what is going on. This is what the questions look like.

OpenStudy (anonymous):

since the degree of the numerator and denominator are the same (both degree 1) the limit is the ratio of the leading coefficients

OpenStudy (kkbrookly):

Okay, so how would I determine if it is converging or diverging?

OpenStudy (kkbrookly):

Ohhhh

OpenStudy (anonymous):

if you get a number, then it converges since the ration of the leading coefficients is \(\frac{1}{5}\) that is your limit

OpenStudy (kkbrookly):

Okay, I understand now. What about the other questions that are not in fraction form?

OpenStudy (anonymous):

second one the degree of the top is smaller than the degree of the bottom limit is 0

OpenStudy (kkbrookly):

Why is it 0 when it is basically the same as the first equation?

OpenStudy (anonymous):

because the degree of the top is smaller than the degree of the bottom replace n by 100 and see what you get

OpenStudy (kkbrookly):

I got .000387..

OpenStudy (kkbrookly):

So that means that the second one isn't converging?

OpenStudy (anonymous):

it does converge it converges to 0 numbers get closer and closer to 0

OpenStudy (kkbrookly):

Alright.. I am starting to understand more. So, for something like number 3 how would I go about finding the limit? Just by plugging in a number to n?

OpenStudy (anonymous):

for number 3, the bigger n gets, the bigger \(\sqrt{25+n}\) gets no limit ( or \(\infty\))

OpenStudy (kkbrookly):

But how would I know that the bigger n gets the bigger the squareroot of 25+n gets?

OpenStudy (anonymous):

not sure what to say other than it is obvious \(\sqrt{100}=10, \sqrt{1,000,000}=1,000\) etc the bigger the number the bigger the square root of the number

OpenStudy (kkbrookly):

I completely understand now. Thank you for being so patient with me!

OpenStudy (anonymous):

yw hope it is clear that the next one \[\sqrt{25+\frac{1}{n}}\] has limit \(5\)

OpenStudy (kkbrookly):

Yes, thank you!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

eat mah poop